# Counting subalgebras of $\mathfrak{o}^{n}$

### 18 Mar 2026 v1

@@ 1. Introduction
1.1. General properties of subring zeta functions
@@ 1.5. Uniformity
We finally mention an important open question concerning the behaviour of the local factors of $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$. Although each $\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is a rational function in $p^{-s}$, it is unknown how this rational expression depends on $p$. We say that $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is uniform if there exists a rational function $W_{n}(X,Y)\in\mathbb{Q}(X,Y)$ such that
@@ 2. Local zeta functions as integrals
In this section, we use the following special notation. For a matrix $A\in\operatorname{Mat}_{m\times n}(R)$, where $R$ is a commutative ring, we denote by $\langle A\rangle$ the $R$-submodule generated by the row vectors of $A$.
@@ 2.1. Parameterization of finite-index submodules
Any submodule $\Lambda\leq\mathfrak{o}^{n}$ is of the form $\Lambda=\langle A\rangle$ for some $A\in\operatorname{Mat}_{n}(\mathfrak{o})$. The submodule $\Lambda$ has finite index if and only if $A$ is nonsingular, i.e. $\det(A)\neq 0$. If $A,B\in\operatorname{Mat}_{n}(\mathfrak{o})$ are nonsingular, then $\langle A\rangle=\langle B\rangle$ if and only if $A=UB$ for some $U\in\operatorname{GL}_{n}(\mathfrak{o})$.
@@
The natural action of $\operatorname{GL}_{n}(R)$ on $R^{n}$ by right multiplication induces a transitive action on the set of $I$-flags. Let $e_{1},\ldots,e_{n}$ be the standard basis of $R^{n}$, and consider the canonical $I$-flag
@@
where $\Delta_{\iota}=\langle e_{j}:j>\iota\rangle$ for $\iota\in I$. If $A\in\operatorname{GL}_{n}(R)$ has row vectors $A_{1},\ldots,A_{n}$, then the image of the canonical flag under $A$ is the flag
@@
Definition 2.2.
A matrix $A\in\operatorname{GL}_{n}(R)$ is called $I$-reduced if it satisfies the following conditions:
@@
    In each row, the leftmost invertible entry is $1$. We call these entries the leading $1$’s.
@@
    All entries above a leading $1$ are zero.
@@
    Setting $i_{0}=0$ and $i_{\ell+1}=n$, for each $j=0,\ldots,\ell$, the leading $1$’s in rows $i_{j}+1,\ldots,i_{j+1}$ occur in strictly increasing column positions.
We denote the set of $I$-reduced matrices by $\mathcal{R}_{I}(R)$.
@@
The following are examples of $I$-reduced matrices for $n=6$ and $I=\{2,4\}$. The leading $1$’s are displayed in bold.
@@
Reduction modulo the maximal ideal of $R$ sends an $I$-reduced matrix over $R$ to an $I$-reduced matrix over the residue field.
@@
Equivalently, every $I$-flag in $R^{n}$ is of the form $\langle A\rangle_{I}$ for a unique $A\in\mathcal{R}_{I}(R)$. We prove this by induction on $n$.
@@
be an $I$-flag. We construct a matrix $A\in\mathcal{R}_{I}(R)$ such that $\langle A\rangle_{I}$ equals this flag.
Choose a basis $\{v_{i_{\ell}+1},\ldots,v_{n}\}\subset R^{n}$ for $\Lambda_{i_{\ell}}$ and apply Gaussian elimination to the matrix with row vectors $v_{i_{\ell}+1},\ldots,v_{n}$. Over a local ring, the only difference from the field case is that one must pivot on invertible entries rather than merely nonzero ones. The resulting matrix has the property that the leftmost invertible entry in each row is $1$, the column positions of these leading $1$’s increase from top to bottom, and all entries above each leading $1$ vanish.
@@
of $R^{n}$ such that, for each $\iota\in I\setminus\{i_{\ell}\}$, the subset $\{v_{\iota+1},\ldots,v_{i_{\ell}},A_{i_{\ell}+1},\ldots,A_{n}\}$ is a basis of $\Lambda_{\iota}$. We may further assume that each $v_{j}$ has zeros in the coordinates corresponding to the leading $1$’s of $A_{i_{\ell}+1},\ldots,A_{n}$. Deleting these coordinates yields vectors $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}\in R^{i_{\ell}}$.
Let $I^{\prime}:=I\setminus\{i_{\ell}\}$. By the induction hypothesis, there exists an $I^{\prime}$-reduced matrix $A^{\prime}\in\operatorname{GL}_{i_{\ell}}(R)$ whose associated $I^{\prime}$-flag coincides with the flag generated by $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}$. Let $A_{1}^{\prime},\ldots,A_{i_{\ell}}^{\prime}$ be the rows of $A^{\prime}$, and lift them to vectors $A_{1},\ldots,A_{i_{\ell}}\in R^{n}$ by reinserting zeros in the deleted coordinates. Then the matrix $A$ with rows $A_{1},\ldots,A_{n}$ is $I$-reduced and satisfies $\langle A\rangle_{I}=\{\Lambda_{\iota}\}$.
We now prove uniqueness. Suppose $A,B\in\mathcal{R}_{I}(R)$ satisfy $\langle A\rangle_{I}=\langle B\rangle_{I}$. Let $A^{\prime\prime}$ (resp. $B^{\prime\prime}$) denote the submatrix consisting of the last $n-i_{\ell}$ rows of $A$ (resp. $B$). Then $\langle A^{\prime\prime}\rangle=\langle B^{\prime\prime}\rangle$, so $A^{\prime\prime}=UB^{\prime\prime}$ for some $U\in\operatorname{GL}_{n-i_{\ell}}(R)$.
Reducing modulo the maximal ideal of $R$, the matrices $\overline{A^{\prime\prime}}$ and $\overline{B^{\prime\prime}}$ are row-equivalent matrices in reduced row echelon form over the residue field, and hence coincide. In particular, the leading $1$’s in $A^{\prime\prime}$ and $B^{\prime\prime}$ occur in the same column positions. Restricting the equality $A^{\prime\prime}=UB^{\prime\prime}$ to these pivot columns yields $U=\operatorname{Id}$, and therefore $A^{\prime\prime}=B^{\prime\prime}$.
Finally, let $A^{\prime}$ (resp. $B^{\prime}$) be obtained from $A$ (resp. $B$) by deleting the last $n-i_{\ell}$ rows and the columns containing the leading $1$’s of those rows. Then $A^{\prime}$ and $B^{\prime}$ are $I^{\prime}$-reduced matrices in $\operatorname{GL}_{i_{\ell}}(R)$ and satisfy $\langle A^{\prime}\rangle_{I^{\prime}}=\langle B^{\prime}\rangle_{I^{\prime}}$. By induction, $A^{\prime}=B^{\prime}$, and hence $A=B$. ∎
@@
The proof yields a bijection between $\mathcal{R}_{I}(R)$ and the set of $I$-flags in $R^{n}$. In particular, for $R=\mathbb{F}_{q}$ we obtain
@@
Let $A\in\mathcal{R}_{I}(\mathfrak{o})$ (or $A\in\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})$). Then $A$ can be written uniquely as $A=B\sigma$ with $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, where $B\sigma$ denotes the matrix obtained from $B$ by permuting its columns according to $\sigma$. Conversely, given $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, the matrix $B\sigma$ is $I$-reduced if and only if $\overline{B}\sigma$ is $I$-reduced, where $\overline{(\cdot)}$ denotes reduction modulo $\mathfrak{m}$.
@@
2.2. Preliminaries on $\pi$-adic integration
@@
Indeed, $\mu^{\prime}\big(\operatorname{L}_{I}(\mathfrak{o})\big)=1$ since $\mathfrak{o}^{N}$ has measure $1$. Fix $B\in\operatorname{L}_{I}(\mathfrak{o})$. Right multiplication by $B$ corresponds to an affine transformation $\varphi:\mathfrak{o}^{N}\to\mathfrak{o}^{N}$ of the form $\varphi(\mathbf{x})=\lambda+T\mathbf{x}$, with $\lambda\in\mathfrak{o}^{N}$ and $T\in\operatorname{Mat}_{N}(\mathfrak{o})$. Since $\varphi$ is bijective, we have $T\in\operatorname{GL}_{N}(\mathfrak{o})$, and hence $\det(T)$ is a unit in $\mathfrak{o}$. Therefore $|\det(T)|=1$, and it follows that
@@ 3. The case #I=1\#I=1
Let $B=\begin{pmatrix}\operatorname{Id}_{\iota}&0\\ \beta^{\prime}&\operatorname{Id}_{n-\iota}\end{pmatrix}\in\operatorname{L}_{I}(\mathfrak{o})$. The condition $\pi x_{\iota}|B_{ij}(B_{ij}-1)\,\forall i,j$ with $\iota+1\leq i\leq n$, $1\leq j\leq\iota$ translates as: each coefficient of $B^{\prime}$ is congruent to 0 or $1$ modulo $\pi x_{\iota}$. The condition $B_{ij}B_{i^{\prime}j}\,\forall i,i^{\prime},j$ with $\iota+1\leq i<i^{\prime}\leq n$, $1\leq j\leq\iota$ translates as: in each column of $B^{\prime}$ there is at most one unit. Hence, condition $\pi x_{\iota}|\mathscr{R}(\beta)$ is equivalent to: $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$ for some very elementary matrix $\epsilon$ and some $B^{\prime\prime}\in\operatorname{Mat}_{(n-\iota)\times\iota}(\mathfrak{o})$. In particular, the integral is 0 if $\delta_{E}=0$. If $\delta_{E}=1$, then $\epsilon$ is completely determined, and we can apply the change of variables $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$. We obtain
@@
The sum to be computed is the number of $I$-reduced matrices over $\mathfrak{o}/\mathfrak{m}$ whose entries are only 0 or $1$ and such that there is at most two non-zero entries in each column. To count these matrices, we first choose the column positions $j_{\iota+1},\ldots,j_{n}$ for the leading 1’s of rows $\iota+1,\ldots,n$. The leading 1’s of the first $\iota$ rows are determined are then completely determined.
@@ 3.1. Application I
We now apply Proposition 3.4. Note that $\binom{2}{1}_{q}=1+q$ and the coefficient of $z$ in $\frac{1}{(1-z)(1-2z)}$ is $3$. Hence
@@ 4.2. The case $n=3$ and $I=\{1,2\}$
      We compute $S_{2,1}$. We show first that the condition $u_{1}\mid a-\chi^{\times}(\alpha)$ can be removed. This is obvious if $u_{1}$ is a unit. If $\pi\mid u_{1}$, then this condition follows from the last condition and the hypothesis $u_{1}u_{4}\mid u_{2}$. Finally, conditions $\pi u_{4}\mid u_{2}$ and $u_{2}\mid\big(a-\chi^{\times}(\alpha)\big)u_{4}$ imply $\alpha-\chi^{\times}(\alpha)=0$, hence $\alpha=0$ or $1$, so the condition $\overline{a}=\alpha$ can be omitted. Therefore,
@@
      If $\alpha=0$ or $1$, then the last integral equals

### 19 Mar 2026 v2

@@ 1. Introduction
$1.$ General properties of subring zeta functions
@@ 1.5. Uniformity
We finally mention an important open question concerning the behaviour of the local factors of $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$. Although each $\zeta_{\mathbb{Z}_{p}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is a rational function in $p^{-s}$, it is unknown how this rational expression depends on pp. We say that $\zeta_{\mathbb{Z}^{n}}^{\scriptscriptstyle{1,<}}(s)$ is uniform if there exists a rational function $W_{n}(X,Y)\in\mathbb{Q}(X,Y)$ such that
@@ 2. Local zeta functions as integrals
In this section, we use the following special notation. For a matrix $A\in\operatorname{Mat}_{m\times n}(R)$, where $R$ is a commutative ring, we denote by $\langle A\rangle$ the $R$-submodule generated by the row vectors of AA.
@@ 2.1. Parameterization of finite-index submodules
Any submodule $\Lambda\leq\mathfrak{o}^{n}$ is of the form $\Lambda=\langle A\rangle$ for some $A\in\operatorname{Mat}_{n}(\mathfrak{o})$. The submodule $\Lambda$ has finite index if and only if AA is nonsingular, i.e. $\det(A)\neq 0$. If $A,B\in\operatorname{Mat}_{n}(\mathfrak{o})$ are nonsingular, then $\langle A\rangle=\langle B\rangle$ if and only if $A=UB$ for some $U\in\operatorname{GL}_{n}(\mathfrak{o})$.
@@
The natural action of $\operatorname{GL}_{n}(R)$ on $R^{n}$ by right multiplication induces a transitive action on the set of II-flags. Let $e_{1},\ldots,e_{n}$ be the standard basis of $R^{n}$, and consider the canonical II-flag
@@
where $\Delta_{\iota}=\langle e_{j}:j>\iota\rangle$ for $\iota\in I$. If $A\in\operatorname{GL}_{n}(R)$ has row vectors $A_{1},\ldots,A_{n}$, then the image of the canonical flag under AA is the flag
@@
Definition $2.$
A matrix $A\in\operatorname{GL}_{n}(R)$ is called II-reduced if it satisfies the following conditions:
@@
    In each row, the leftmost invertible entry is 11. We call these entries the leading 11’s.
@@
    All entries above a leading 11 are zero.
@@
    Setting $i_{0}=0$ and $i_{\ell+1}=n$, for each $j=0,\ldots,\ell$, the leading 11’s in rows $i_{j}+1,\ldots,i_{j+1}$ occur in strictly increasing column positions.
We denote the set of II-reduced matrices by $\mathcal{R}_{I}(R)$.
@@
The following are examples of II-reduced matrices for $n=6$ and $I=\{2,4\}$. The leading 11’s are displayed in bold.
@@
Reduction modulo the maximal ideal of $R$ sends an II-reduced matrix over $R$ to an II-reduced matrix over the residue field.
@@
Equivalently, every II-flag in $R^{n}$ is of the form $\langle A\rangle_{I}$ for a unique $A\in\mathcal{R}_{I}(R)$. We prove this by induction on $n$.
@@
be an II-flag. We construct a matrix $A\in\mathcal{R}_{I}(R)$ such that $\langle A\rangle_{I}$ equals this flag.
Choose a basis $\{v_{i_{\ell}+1},\ldots,v_{n}\}\subset R^{n}$ for $\Lambda_{i_{\ell}}$ and apply Gaussian elimination to the matrix with row vectors $v_{i_{\ell}+1},\ldots,v_{n}$. Over a local ring, the only difference from the field case is that one must pivot on invertible entries rather than merely nonzero ones. The resulting matrix has the property that the leftmost invertible entry in each row is 11, the column positions of these leading 11’s increase from top to bottom, and all entries above each leading 11 vanish.
@@
of $R^{n}$ such that, for each $\iota\in I\setminus\{i_{\ell}\}$, the subset $\{v_{\iota+1},\ldots,v_{i_{\ell}},A_{i_{\ell}+1},\ldots,A_{n}\}$ is a basis of $\Lambda_{\iota}$. We may further assume that each $v_{j}$ has zeros in the coordinates corresponding to the leading 11’s of $A_{i_{\ell}+1},\ldots,A_{n}$. Deleting these coordinates yields vectors $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}\in R^{i_{\ell}}$.
Let $I^{\prime}:=I\setminus\{i_{\ell}\}$. By the induction hypothesis, there exists an $I^{\prime}$-reduced matrix $A^{\prime}\in\operatorname{GL}_{i_{\ell}}(R)$ whose associated $I^{\prime}$-flag coincides with the flag generated by $v_{1}^{\prime},\ldots,v_{i_{\ell}}^{\prime}$. Let $A_{1}^{\prime},\ldots,A_{i_{\ell}}^{\prime}$ be the rows of $A^{\prime}$, and lift them to vectors $A_{1},\ldots,A_{i_{\ell}}\in R^{n}$ by reinserting zeros in the deleted coordinates. Then the matrix AA with rows $A_{1},\ldots,A_{n}$ is II-reduced and satisfies $\langle A\rangle_{I}=\{\Lambda_{\iota}\}$.
We now prove uniqueness. Suppose $A,B\in\mathcal{R}_{I}(R)$ satisfy $\langle A\rangle_{I}=\langle B\rangle_{I}$. Let $A^{\prime\prime}$ (resp. $B^{\prime\prime}$) denote the submatrix consisting of the last $n-i_{\ell}$ rows of AA (resp. $B$). Then $\langle A^{\prime\prime}\rangle=\langle B^{\prime\prime}\rangle$, so $A^{\prime\prime}=UB^{\prime\prime}$ for some $U\in\operatorname{GL}_{n-i_{\ell}}(R)$.
Reducing modulo the maximal ideal of $R$, the matrices $\overline{A^{\prime\prime}}$ and $\overline{B^{\prime\prime}}$ are row-equivalent matrices in reduced row echelon form over the residue field, and hence coincide. In particular, the leading 11’s in $A^{\prime\prime}$ and $B^{\prime\prime}$ occur in the same column positions. Restricting the equality $A^{\prime\prime}=UB^{\prime\prime}$ to these pivot columns yields $U=\operatorname{Id}$, and therefore $A^{\prime\prime}=B^{\prime\prime}$.
Finally, let $A^{\prime}$ (resp. $B^{\prime}$) be obtained from AA (resp. $B$) by deleting the last $n-i_{\ell}$ rows and the columns containing the leading 11’s of those rows. Then $A^{\prime}$ and $B^{\prime}$ are $I^{\prime}$-reduced matrices in $\operatorname{GL}_{i_{\ell}}(R)$ and satisfy $\langle A^{\prime}\rangle_{I^{\prime}}=\langle B^{\prime}\rangle_{I^{\prime}}$. By induction, $A^{\prime}=B^{\prime}$, and hence $A=B$. ∎
@@
The proof yields a bijection between $\mathcal{R}_{I}(R)$ and the set of II-flags in $R^{n}$. In particular, for $R=\mathbb{F}_{q}$ we obtain
@@
Let $A\in\mathcal{R}_{I}(\mathfrak{o})$ (or $A\in\mathcal{R}_{I}(\mathfrak{o}/\mathfrak{m})$). Then AA can be written uniquely as $A=B\sigma$ with $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, where $B\sigma$ denotes the matrix obtained from $B$ by permuting its columns according to $\sigma$. Conversely, given $B\in\operatorname{L}_{I}(\mathfrak{o})$ and $\sigma\in\mathbb{S}_{n}$, the matrix $B\sigma$ is II-reduced if and only if $\overline{B}\sigma$ is II-reduced, where $\overline{(\cdot)}$ denotes reduction modulo $\mathfrak{m}$.
@@
$2.$ Preliminaries on $\pi$-adic integration
@@
Indeed, $\mu^{\prime}\big(\operatorname{L}_{I}(\mathfrak{o})\big)=1$ since $\mathfrak{o}^{N}$ has measure 11. Fix $B\in\operatorname{L}_{I}(\mathfrak{o})$. Right multiplication by $B$ corresponds to an affine transformation $\varphi:\mathfrak{o}^{N}\to\mathfrak{o}^{N}$ of the form $\varphi(\mathbf{x})=\lambda+T\mathbf{x}$, with $\lambda\in\mathfrak{o}^{N}$ and $T\in\operatorname{Mat}_{N}(\mathfrak{o})$. Since $\varphi$ is bijective, we have $T\in\operatorname{GL}_{N}(\mathfrak{o})$, and hence $\det(T)$ is a unit in $\mathfrak{o}$. Therefore $|\det(T)|=1$, and it follows that
@@ 3. The case #I=1\#I=1
Let $B=\begin{pmatrix}\operatorname{Id}_{\iota}&0\\ \beta^{\prime}&\operatorname{Id}_{n-\iota}\end{pmatrix}\in\operatorname{L}_{I}(\mathfrak{o})$. The condition $\pi x_{\iota}|B_{ij}(B_{ij}-1)\,\forall i,j$ with $\iota+1\leq i\leq n$, $1\leq j\leq\iota$ translates as: each coefficient of $B^{\prime}$ is congruent to 0 or 11 modulo $\pi x_{\iota}$. The condition $B_{ij}B_{i^{\prime}j}\,\forall i,i^{\prime},j$ with $\iota+1\leq i<i^{\prime}\leq n$, $1\leq j\leq\iota$ translates as: in each column of $B^{\prime}$ there is at most one unit. Hence, condition $\pi x_{\iota}|\mathscr{R}(\beta)$ is equivalent to: $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$ for some very elementary matrix $\epsilon$ and some $B^{\prime\prime}\in\operatorname{Mat}_{(n-\iota)\times\iota}(\mathfrak{o})$. In particular, the integral is 0 if $\delta_{E}=0$. If $\delta_{E}=1$, then $\epsilon$ is completely determined, and we can apply the change of variables $B^{\prime}=\epsilon+\pi x_{\iota}B^{\prime\prime}$. We obtain
@@
The sum to be computed is the number of II-reduced matrices over $\mathfrak{o}/\mathfrak{m}$ whose entries are only 0 or 11 and such that there is at most two non-zero entries in each column. To count these matrices, we first choose the column positions $j_{\iota+1},\ldots,j_{n}$ for the leading 1’s of rows $\iota+1,\ldots,n$. The leading 1’s of the first $\iota$ rows are determined are then completely determined.
@@ 3.1. Application I
We now apply Proposition 3.4. Note that $\binom{2}{1}_{q}=1+q$ and the coefficient of $z$ in $\frac{1}{(1-z)(1-2z)}$ is 33. Hence
@@ 4.2. The case $n=3$ and $I=\{1,2\}$
      We compute $S_{2,1}$. We show first that the condition $u_{1}\mid a-\chi^{\times}(\alpha)$ can be removed. This is obvious if $u_{1}$ is a unit. If $\pi\mid u_{1}$, then this condition follows from the last condition and the hypothesis $u_{1}u_{4}\mid u_{2}$. Finally, conditions $\pi u_{4}\mid u_{2}$ and $u_{2}\mid\big(a-\chi^{\times}(\alpha)\big)u_{4}$ imply $\alpha-\chi^{\times}(\alpha)=0$, hence $\alpha=0$ or 11, so the condition $\overline{a}=\alpha$ can be omitted. Therefore,
@@
      If $\alpha=0$ or 11, then the last integral equals